# STONet: A novel neural operator for modeling solute transport in micro-cracked reservoirs

### v1

@@ Abstract
In this work, we develop a novel neural operator, the Solute Transport Operator Network (STONet), to efficiently model contaminant transport in micro-cracked reservoirs. The model combines different networks to encode heterogeneous properties effectively. By predicting the concentration rate, we are able to accurately model the transport process. Numerical experiments demonstrate that our neural operator approach achieves accuracy comparable to that of the finite element method. The previously introduced Enriched DeepONet architecture has been revised, motivated by the architecture of the popular multi-head attention of transformers, to improve its performance without increasing the compute cost. The computational efficiency of the proposed model enables rapid and accurate predictions of solute transport, facilitating the optimization of reservoir management strategies and the assessment of environmental impacts. The data and code for the paper will be published at https://github.com/ehsanhaghighat/STONet.
Keywords: Neural Operators; Porous Media; Solute Transport.
@@ 1.   Introduction
The depletion of freshwater resources is a pressing global challenge, particularly in regions facing severe droughts leading to the rapid exhaustion of groundwater reserves. A significant factor contributing to water quality degradation in underground aquifers is the intrusion of seawater: the higher density of saline water facilitates its rapid dispersion and mixing within freshwater aquifers, leading to the groundwater contamination [5, 6]. Assessing the risk of seawater intrusion and developing mitigating strategies requires quantitative modeling of coupled flow and solute transport in porous media [1, 28]. These assessments are further complicated by the common occurrence of fractures in the subsurface, which can significantly alter the flow: typically, fractures exhibit higher permeability than the surrounding domain, thus profoundly modulating groundwater flow and transport [12, 23, 30, 7, 8, 16].
@@
Over the past few years, there has been an explosive increase in the development and application of deep learning (DL) approaches, partly as a result of data availability and computing power [20]. Recent advances in deep learning approaches have pushed engineers and scientists to leverage ML frameworks for solving classical engineering problems. A recent class of DL methods, namely, Phyics-Informed Neural Networks (PINNs), have received increased attention for solving forward and inverse problems and for building surrogate models with lesser data requirements [27, 13, 4]. PINNs leverage physical principles and incorporate them into the optimization process, enabling the network to learn the underlying physics of the problem. The applications of this approach extend to fluid mechanics, solid mechanics, heat transfer, and flow in porous media, among others [9, 3, 25, 10, 2, 11]. A recent architecture, namely Neural Operators, provides an efficient framework for data-driven and physics-informed surrogate modeling [22, 21, 33, 34, 19]. Neural Operators are a class of neural networks that operate on functions rather than vectors, enabling them to capture the relationships between input and output functions. By leveraging the power of Neural Operators, it is possible to construct surrogate models that are both data-efficient and accurate. This makes Neural Operators well-suited for problems where experimental data is limited or computationally expensive to obtain. Once trained, neural operators can be used to perform inference efficiently.
@@
In this study, we developed a novel neural operator specifically designed for modeling density-driven flow in fractured porous media. The neural operator leverages the power of deep learning to capture the complex relationships between the equivalent permeability tensor, which is a result of variations in fracture orientation and fracture density, and the pressure gradient, and output the temporally varying concentration field. The new architecture, as detailed in section 3.2, revises the previously introduced En-DeepONet [11] to achieve higher accuracy at the same computational cost.
We generated a comprehensive dataset obtained using high-fidelity finite element simulations to train and validate the neural operator. This dataset serves as a reliable benchmark for evaluating the performance of the proposed neural operator. We applied the developed framework to predict flow patterns in porous media under heterogeneous conditions, demonstrating its ability to handle complex geological scenarios.
@@ 2.   Governing Equations
Here, the first term in eq. (7) denotes the rate of change in the mass of the solute component, while the second and third terms denote the advective and dispersive transport mechanisms of the solute component, respectively. $D_{m}$ refers to the molecular diffusion coefficient associated with the matrix, $\tau$ represents the tortuosity of the porous medium, $\mathbf{I}$ is the identity tensor and $\alpha_{L}$ and $\alpha_{T}$ denotes the longitudinal and transverse dispersivities, respectively. The solution to the aforementioned equations can be achieved through the finite element method, as detailed in [15].
@@ 3.1 Enriched DeepONet
where $x\in X$ and $u\in U$ are a query point in the solution domain $\Omega$ and a problem parameter set, respectively. Here, $\boldsymbol{\theta}_{\alpha}$ represents the set of parameters of each network, and $\epsilon_{B},\epsilon_{T}$ are encoded outputs of branch and trunk networks, respectively. ,$\odot$,$\oplus$,$\ominus$ , ⊕ , ⊖ denotes elementwise multiplication, addition, and subtraction of branch and trunk encodings, respectively. $R$ is a fusion network, known as the root network, which decodes the final outputs.
Figure 1: Network architecture. (a) En-DeepONet neural architecture. (b) The revised En-DeepONet architecture, namely STONet. STONet resembles the multi-head attention mechanism of transformer architecture and applies the multiplication, addition, and subtraction operations on different layers.
@@ 3.2 STONet
Here, we present an extension of the En-DeepONet architecture that resembles the multi-head attention mechanism of transformer networks [31]. The network architecture is depicted in fig. 1(b). The architecture consists of an encoding branch and a trunk network. The output of the branch network is then combined with the output of the trunk network, using elementwise operations such as multiplication or addition, and passed to the new attention block. One may add multiple attention blocks. The final output is then passed on to the output root network. The network architecture is expressed mathematically as
@@ 4.   Results and Discussion
4.1 Problem Description
A schematic of the problem considered in this study of a solute through a micro-cracked reservoir is shown in fig. 2. The domain under consideration has dimensions of $70\times 50$ cm. The solution is injected from the left side and transported through the domain, driven by a pressure difference between the left and right boundaries. The porous reservoir is assumed to be confined between two impervious layers and contains randomly distributed micro-cracks of varying density and orientations. The micro-fracture orientations are sampled randomly from a normal distribution with varying mean values but fixed standard deviation, i.e., $\theta\sim\mathcal{N}(\mu_{\theta},\sigma_{\theta}=15^{\circ})$, where $\mu_{\theta}$ is sampled uniformly as $\mu_{\theta}\sim\mathcal{U}(-60^{\circ},60^{\circ})$. The representative elementary volume (REV) dimensions for assessing the permeability field (using eq. (5)) are assumed $10\times 10$ cm. Fracture density, i.e., the number of cracks per REV, is sampled from a Poisson’s distribution as $\Sigma\sim\mathcal{P}(\lambda)$, where $\lambda$ is sampled uniformly as $\lambda\sim\mathcal{U}(30,70)$. Fracture length and aperture are sampled from log-normal distributions $\log(l_{c})\sim\mathcal{N}(\mu_{l_{c}}=0.05,\sigma_{\l_{c}}=0.0575$ and $\log(a)\sim\mathcal{N}(\mu_{a}=1.14\times 10^{-4},\sigma_{a}=1.15\times 10^{-4}\times 1.5)$, respectively. The pressure on the left side is kept fixed, while the pressure on the right side is also randomly perturbed from a normal distribution, i.e., $p_{right}=\mathcal{U}(4976,4996)+9792.34y$. fig. 3 depicts a sample medium with micro-fractures.
@@ 4.1 Problem Description
Figure 4 depicts a few snapshots of the solute concentration for different micro-crack distributions. The equivalent permeability field $k_{x}$ ($k_{11}$) and $k_{y}$ ($k_{22}$) are plotted in the left two columns, while the right two columns highlight concentration and the rate of change of concentration at the last time step (i.e., $t=36$ hr). The changes in the concentration patterns are due to different fracture orientations and pressure gradients.
@@ 4.2 Sampling Strategy
Since concentration remains near zero for a large portion of the domain (as shown in fig. 4), to reduce the batch size and computational demand, we sub-sample 1,500 random nodes using an importance sampling strategy. Out of these, 1,000 points were selected based on concentration density, while the remaining 500 nodes were uniformly distributed over the domain.
@@ 4.3 Training Performance
Figure 5: The performance of the new STONet architecture versus the old En-DeepONet architecture. The x-axis shows the total number of parameters, and the y-axis presents the average loss values of the last 20 epochs for training performed for 500 epochs. Each subplot show the variations with respect to the width of the network, number of branch and trunk layers, number of root layers, and number of attention blocks, respectively.

### v2

# STONet: A neural operator for modeling solute transport in micro-cracked reservoirs

@@ Abstract
In this work, we introduce a novel neural operator, the Solute Transport Operator Network (STONet), to efficiently model contaminant transport in micro-cracked porous media. STONet’s model architecture is specifically designed for this problem and uniquely integrates an enriched DeepONet structure with a transformer-based multi-head attention mechanism, enhancing performance without incurring additional computational overhead compared to existing neural operators. The model combines different networks to encode heterogeneous properties effectively and predict the rate of change of the concentration field to accurately model the transport process. The training data is obtained using finite element (FEM) simulations by random sampling of micro-fracture distributions and applied pressure boundary conditions, which capture diverse scenarios of fracture densities, orientations, apertures, lengths, and balance of pressure-driven to density-driven flow. Our numerical experiments demonstrate that, once trained, STONet achieves accurate predictions, with relative errors typically below 1% compared with FEM simulations while reducing runtime by approximately two orders of magnitude. This type of computational efficiency facilitates building digital twins for rapid assessment of subsurface contamination risks and optimization of environmental remediation strategies. The data and code for the paper will be published at https://github.com/ehsanhaghighat/STONet.
Keywords: Machine Learning; Neural Operators; Fractured Porous Media; Solute Transport.
@@ 1.   Introduction
The depletion of freshwater resources is a pressing global challenge, particularly in regions facing severe droughts leading to the rapid exhaustion of groundwater reserves. A significant factor contributing to water quality degradation in underground aquifers is the intrusion of seawater: the higher density of saline water facilitates its rapid dispersion and mixing within freshwater aquifers, leading to the groundwater contamination [6, 7]. Assessing the risk of seawater intrusion and developing mitigating strategies requires quantitative modeling of coupled flow and solute transport in porous media [1, 41]. These assessments are further complicated by the common occurrence of fractures in the subsurface, which can significantly alter the flow: typically, fractures exhibit higher permeability than the surrounding domain, thus profoundly modulating groundwater flow and transport [22, 34, 46, 8, 10, 19]. Other factors that can affect the solute transport problem include, but not limited to, transport under partially saturated conditions [21, 57, 42] and as well as other environmental and mechanical conditions such as erosion [39, 35, 43].
@@
Numerical modeling techniques, also known as forward models, such as the finite element method (FEM) have traditionally been employed to simulate flow and transport in porous media [44]. Forward models rely on an accurate understanding of model parameters, which are mostly unknown for subsurface applications except at sparse observation or injection/production wells. Therefore, repeated simulations are often performed to find model parameters while matching the data at wells. Although powerful, each forward simulation is computationally intensive, particularly for realistic, three-dimensional simulations, where computational runtimes can extend to several hours or even days for a single scenario. This significant computational demand severely limits their applicability for real-time analyses and identification or optimization tasks, where numerous simulations are necessary to explore parameter spaces or identify optimal reservoir management strategies. In contrast, machine learning (ML) techniques, particularly neural operators, offer highly efficient inference capabilities once trained, with prediction times typically reduced to seconds or less [33, 3, 30]. Furthermore, ML models inherently provide analytical differentiation, a feature invaluable for optimization and sensitivity analysis. As a result, ML-based surrogate models have increasingly become an attractive and practical solution for rapid assessment and optimization of complex subsurface and groundwater contamination problems.
@@
Over the past few years, there has been an explosive increase in the development and application of deep learning (DL) approaches, partly as a result of data availability and computing power [31]. Recent advances in deep learning approaches have pushed engineers and scientists to leverage ML frameworks for solving classical engineering problems. A recent class of DL methods, namely, Phyics-Informed Neural Networks (PINNs), have received increased attention for solving forward and inverse problems and for building surrogate models with lesser data requirements [40, 23, 5]. PINNs leverage physical principles and incorporate them into the optimization process, enabling the network to learn the underlying physics of the problem. The applications of this approach extend to fluid mechanics, solid mechanics, heat transfer, and flow in porous media, among others [20, 24, 11, 14, 4, 37, 15, 12, 2, 52, 55, 53]. A recent architecture, namely Neural Operators, provides an efficient framework for data-driven and physics-informed surrogate modeling [33, 32, 49, 49, 50, 13, 29, 16]. Neural Operators are a class of neural networks that operate on functions rather than vectors, enabling them to capture the relationships between input and output functions. By leveraging the power of Neural Operators, it is possible to construct surrogate models that are both data-efficient and accurate. This makes Neural Operators well-suited for problems where experimental data is limited or computationally expensive to obtain. Once trained, neural operators can be used to perform inference efficiently. Neural operators have recently been used to model transport in porous media [51, 9, 17].
@@
In this study, we developed a neural operator specifically designed for modeling density-driven flow in fractured porous media. The neural operator leverages the power of deep learning to capture the complex relationships between the equivalent permeability tensor, which is a result of variations in fracture orientation and fracture density, and the pressure gradient, and outputs the spatio-temporally varying concentration field. The new architecture, as detailed in section 3.2, revises the previously introduced En-DeepONet [16] to achieve higher accuracy at the same computational cost.
A key contribution of STONet is its careful selection and encoding of input and output features tailored for the flow and solute transport problem. The input features include the equivalent permeability tensor which encodes both intrinsic permeability as well as fracture statistics (orientation, density, length, aperture), and pressure boundary conditions. The output is the rate of change of concentration, which enables flexible and stable auto-regressive prediction of the concentration field over time. Additionally, STONet introduces a novel attention block with residual connections that operates on the encoded features from the branch and trunk networks. The residual connections ensure stable training and facilitate the propagation of important information from the branch network, which encodes the spatially varying physical properties, throughout the network. This design leads to improved accuracy and generalization compared to other DeepONet architectures.
We generated a dataset obtained using high-fidelity finite element simulations to train and validate the neural operator. This dataset serves as a benchmark for evaluating the performance of the proposed neural operator. We applied the developed framework to predict flow patterns in porous media under heterogeneous conditions, demonstrating its ability to handle complex geological scenarios.
@@ 2.   Governing Equations
Here, the first term in eq. (7) denotes the rate of change in the mass of the solute component, while the second and third terms denote the advective and dispersive transport mechanisms of the solute component, respectively. $D_{m}$ refers to the molecular diffusion coefficient associated with the matrix, $\tau$ represents the tortuosity of the porous medium, $\mathbf{I}$ is the identity tensor, and $\alpha_{L}$ and $\alpha_{T}$ denote the longitudinal and transverse dispersivities, respectively. The solution to the aforementioned equations can be achieved through the finite element method, as detailed in [26].
@@ 3.1 Enriched DeepONet
where $x\in X$ and $u\in U$ are a query point in the solution domain $\Omega$ and a problem parameter set, respectively. Here, $\boldsymbol{\theta}_{\alpha}$ represents the set of parameters of each network, and $\epsilon_{B},\epsilon_{T}$ are encoded outputs of the branch and trunk networks, respectively. ,$\odot$,$\oplus$,$\ominus$ , ⊕ , ⊖ denote elementwise multiplication, addition, and subtraction of branch and trunk encodings, respectively. $R$ is a fusion network, known as the root network, which decodes the final outputs.
Figure 1: Network architecture. (a) En-DeepONet neural architecture [16]. (b) The revised En-DeepONet architecture, namely STONet. STONet resembles the multi-head attention mechanism of transformer architecture [47] with residual connections [18] and applies the multiplication, addition, and subtraction operations on different layers.
@@ 3.2 STONet
Here, we present an extension of the En-DeepONet architecture that resembles the multi-head attention mechanism of transformer networks [47] with residual connections [18]. The network architecture is depicted in fig. 1(b). The architecture consists of an encoding branch and a trunk network. The output of the branch network is then combined with the output of the trunk network, using elementwise operations such as multiplication or addition, and passed to the new attention block. One may add multiple attention blocks. The final output is then passed on to the output root network. The network architecture is expressed mathematically as
@@
Regarding the choice of input and output features, we performed many experiments to find the best-performing architecture, but we avoided reporting all cases here. Those experiments include:
  • Predicting $c$ instead of $\dot{c}$. This is not a good choice because it imposes a fixed-time stepping on the auto-regressive updates, therefore less generalizable. However, we still tested the architecture but we did not observe any improvements on this dataset.
  • Adding velocity field $\mathbf{v}_{m}$ as input and output features. Based on the transport eq. 7, it is clear $\mathbf{v}_{m}$ plays an important role in the transport of the contaminant in the domain. Adding this, however, did not improve the results. Note that $\mathbf{v}_{m}$ is strongly correlated with other features already included, including, equivalent permeability field and boundary condition ($p_{\text{left}}-p_{\text{right}}$). We suspect that the reason it did not improve the results here is because of the simplicity of the domain and the overall distribution of data. However, for more generic cases, the velocity field (or pressure gradient) should also be added as input and output features, possibly as a separate network, to achieve more generalization. Worth noting that adding $\mathbf{v}_{m}$ increased the training time and memory requirements.
@@ 4.   Results and Discussion
4.1 Problem Description and Parameter Specification
A schematic of the problem considered in this study of a solute through a micro-cracked reservoir is shown in fig. 2. The domain under consideration has dimensions of $70\times 50$ cm. The deterministic parameters of the problem are given in Table 1. The solution is injected from the left side and transported through the domain, driven by a pressure difference between the left and right boundaries. The porous reservoir is assumed to be confined between two impervious layers and contains randomly distributed micro-cracks of varying density and orientations. The micro-fracture orientations are sampled randomly from a normal distribution with varying mean values but fixed standard deviation, i.e., $\theta\sim\mathcal{N}(\mu_{\theta},\sigma_{\theta}=15^{\circ})$, where $\mu_{\theta}$ is sampled uniformly as $\mu_{\theta}\sim\mathcal{U}(-60^{\circ},60^{\circ})$. The representative elementary volume (REV) dimensions for assessing the permeability field (using eq. (5)) are assumed $10\times 10$ cm. Fracture density, i.e., the number of cracks per REV, is sampled from a Poisson’s distribution as $\Sigma\sim\mathcal{P}(\lambda)$, where $\lambda$ is sampled uniformly as $\lambda\sim\mathcal{U}(30,70)$. Fracture length and aperture are sampled from log-normal distributions $\log(l_{c})\sim\mathcal{N}(\mu_{l_{c}}=0.05,\sigma_{\l_{c}}=0.0575$ and $\log(a)\sim\mathcal{N}(\mu_{a}=1.14\times 10^{-4},\sigma_{a}=1.15\times 10^{-4}\times 1.5)$, respectively. The pressure on the left side is kept fixed, while the pressure on the right side is also randomly perturbed from a normal distribution, i.e., $p_{right}=\mathcal{U}(4976,4996)+9792.34y$. Figure 3 depicts a sample medium with micro-fractures. The sampling procedure for finding stochastic parameters is described in Algorithm 1.
@@ 4.1 Problem Description and Parameter Specification
Table 1: Deterministic parameters.
Parameter Symbol Value Units
Gravitational acceleration g 9.81 ($m/s^{2}$)
Water density $\rho_{0}$ 998.2 (kg/$m^{3}$)
Brine density $\rho_{s}$ 1002 ($kg/m^{3}$)
Viscosity $\mu$ $1.002\times 10^{-3}$ (Pa.$s$)
Porosity $\phi$ 0.38
Diffusion coefficient $D_{m}$ $1.61\times 10^{-9}$ ($m^{2}$/s)
Intrinsic permeability $k_{r}$ $5.7\times 10^{-11}$ ($m^{2}$)
Longitudinal dispersivity $\alpha_{l}$ $1\times 10^{-3}$ ($m$)
Transverse dispersivity $\alpha_{t}$ $2\times 10^{-4}$ ($m$)
Tortuosity $\tau$ 1
Algorithm 1 Sampling steps for stochastic parameters
1:$N$-samples $\leftarrow$ Total number of samples (500).
2:$N$-quadrature $\leftarrow$ Total number of quadrature points inside the domain.
3:for $k\leftarrow 1$ to $N$-samples do
4:     Sample global parameters (fixed throughout the domain)
5:     $\mu_{\theta}\sim\mathcal{U}(-60^{\circ},60^{\circ})\leftarrow$ Sample $\mu_{\theta}$ uniformly between $-60^{\circ}$ and $60^{\circ}$ for the distribution of micro-fractures within the domain.
6:     $\lambda\sim\mathcal{U}(30,70)\leftarrow$ The global Poisson distribution parameter of fracture density, sampled uniformly between 30 and 70.
7:     Sample local parameters (at each integration point)
8:     for $i\leftarrow 1$ to $N$-quadratures do
9:         $\theta^{i}\sim\mathcal{N}(\mu_{\theta},\sigma_{\theta}=15^{\circ})\leftarrow$ Sample $\theta$ at every integration point inside the discretized Finite Element mesh.
10:         $\Sigma^{i}\sim\mathcal{P}(\lambda)\leftarrow$ Sample fracture density at each integration point using the global value for $\lambda$.
11:         $\log(l_{c}^{i})\sim\mathcal{N}(\mu_{l_{c}}=0.05,\sigma_{l_{c}}=0.0575\leftarrow$ Sample fracture length per integration point.
12:         $\log(a^{i})\sim\mathcal{N}(\mu_{a}=1.14\times 10^{-4},\sigma_{a}=1.15\times 10^{-4}\times 1.5)\leftarrow$ Sample fracture aperture per integration point.
13:         Eval equivalent permeability tensor
14:         $\Omega_{i}\leftarrow$ The Set of all quadrature points within REV defined around integration point $i$
15:         $\mathbf{k}_{r}\leftarrow$ Intrinsic permeability of pore structure.
16:         $\mathbf{M}_{i}\leftarrow$ Compute conversion matrix over $\Omega_{i}$, defined in eq. 6.
17:         $\mathbf{k}_{m}\leftarrow$ Compute equivalent permeability tensor for integration point $i$ over $\Omega_{i}$.
18:     end for
19:     $c\leftarrow$ Solve for concentration for each sample and at each time step.
20:     $\dot{c}\leftarrow$ Eval $\dot{c}$ for each sample and at each time step using backward Euler.
21:     Store data for sample $k$ and move to next simulation.
22:end for
Figure 4 depicts a few snapshots of the solute concentration for different micro-crack distributions. The equivalent permeability field $k_{x}$ ($k_{11}$) and $k_{y}$ ($k_{22}$) are plotted in the left two columns, while the right two columns highlight concentration and the rate of change of concentration at the last time step (i.e., $t=36$ hr). The changes in the concentration patterns are due to different fracture orientations and pressure gradients (or velocity field $\mathbf{v}_{m}$), as shown in fig. 5.
@@
Figure 5: Sample pressure and velocity field for different fracture patterns. Each row represents a different realization, corresponding to each sample in fig. 4. The first column shows the absolute pressure field, the second column shows the relative pressure field (by subtracting the background pressure), and the third and fourth columns show the $x$ and $y$ components of the velocity field $\mathbf{v}_{m}$, respectively, at the last time step (i.e., $t=36$ h).
@@ 4.2 Sampling Strategy
Since concentration remains near zero for a large portion of the domain (as shown in fig. 4), to reduce the batch size and computational demand, we sub-sample 1,500 random nodes using an importance sampling strategy. Out of these, 1,000 points were selected based on concentration density, while the remaining 500 nodes were uniformly distributed over the domain. Therefore, sampling points are random in space and over time.
@@ 4.3 Training Performance
Figure 6: The training performance of the new STONet architecture versus the old En-DeepONet architecture. The x-axis shows the total number of parameters, and the y-axis presents the average loss values of the last 20 epochs for training performed for 500 epochs. Each subplot show the variations with respect to the width of the network, number of branch and trunk layers, number of root layers, and number of attention blocks, respectively.
@@ 5.   Conclusions
While STONet demonstrates promising results for modeling solute transport in micro-cracked reservoirs, several limitations remain. The current study is restricted to a single domain geometry and accounts for limited variations in permeability but doesn’t account for the presence of distinct geological layers. Additionally, the model assumes a fixed injection location for the contaminant and does not consider scenarios where the contaminant flow originates from multiple or varying locations. The study also does not account for different pressure or flux boundary conditions, which are critical for capturing diverse real-world scenarios. Furthermore, the study is limited to a two-dimensional domain, whereas real-world applications often require three-dimensional modeling to capture the full complexity of subsurface transport. Considering all these choices may require hundreds of thousands of training samples, which is out of the scope of the current study. These limitations highlight the need for further research and development to extend the applicability of STONet to more generic and realistic scenarios.